# Rethinking Open-Vocabulary Segmentation of Radiance Fields in 3D Space

### v1

@@ Abstract
Understanding the 3D semantics of a scene is a fundamental problem for various scenarios such as embodied agents. While NeRFs and 3DGS excel at novel-view synthesis, previous methods for understanding their semantics have been limited to incomplete 3D understanding: their segmentation results are 2D masks and their supervision is anchored at 2D pixels. This paper revisits the problem set to pursue a better 3D understanding of a scene modeled by NeRFs and 3DGS as follows. 1) We directly supervise the 3D points to train the language embedding field. It achieves state-of-the-art accuracy without relying on multi-scale language embeddings. 2) We transfer the pre-trained language field to 3DGS, achieving the first real-time rendering speed without sacrificing training time or accuracy. 3) We introduce a 3D querying and evaluation protocol for assessing the reconstructed geometry and semantics together. Code, checkpoints, and annotations will be available online.
Project page: https://hyunji12.github.io/Open3DRF
@@ 1 Introduction
Semantically understanding 3D space is important for various computer vision tasks. For instance, it is crucial to segment 3D objects accurately for robot manipulation (Rashid et al. 2023; Zheng et al. 2024). Recently, several works have focused on understanding 3D scenes represented by radiance fields such as Neural Radiance Fields (NeRFs) and 3D Gaussian Splatting (3DGS). LERF (Kerr et al. 2023) locates objects with open-vocabulary querying in a given viewpoint by introducing a language embedding field. The language field is supervised by CLIP embeddings from the multi-scale patches to capture various sizes of objects. Subsequent works (Zhang, Li, and Ahuja 2024; Qin et al. 2023) incorporate SAM masks (Kirillov et al. 2023) to supervise the language field for clear segmentation boundaries.
@@
We revisit the 3D understanding of NeRFs and 3DGS in four aspects: problem setting, supervision, embeddings, and evaluation. The problem setting of previous works leads to limited 3D understanding as they merely produce 2D masks for given camera viewpoints rather than 3D volume. On the other hand, we set the problem to segment 3D volume regarding the semantics of 3D points.
@@
The previous methods encourage the rendered embeddings, rather than the embeddings in 3D space, to match the ground truth. On top of that, multi-scale language embeddings require finding the optimal scale for a given viewpoint and a query text, leading to view inconsistency. By contrast, we encourage 3D points to learn language embeddings following our revisited problem setting. Although the ground truth is still language embeddings of 2D images, changing the dimension for computing the loss greatly improves 3D understanding as shown in Figure 1. Furthermore, we remove multi-scale embeddings by encoding masked objects, achieving consistent 2D understanding between viewpoints.
@@
In addition, previous works on understanding 3DGS explicitly add language embeddings for each Gaussian and jointly train them (Zhang, Li, and Ahuja 2024; Qin et al. 2023). However, directly appending a 512-dimensional language embedding to each Gaussian and rendering them exceeds the maximum GPU shared memory by 10. Hence, they either 1) compress language embeddings to low-dimensional features, significantly sacrificing the accuracy (Shi et al. 2023; Qin et al. 2023), or 2) utilize GPU global memory which is slow for training (Zhou et al. 2023; Ye and Kanazawa 2023). To overcome these drawbacks, we transfer our pre-trained language field into 3DGS to enable the first real-time rendering of the accurate language field.
@@
Our final aspect is evaluation. Existing works only report their accuracy in mIoU for given viewpoints as in 2D segmentation, or measure mIoU on the ground truth point clouds instead of reconstructed geometry (Engelmann et al. 2024). However, 3D understanding should know the correct 3D volume of a target semantics. Therefore, the evaluation method should assess semantics and the reconstructed geometry together. Accordingly, we propose to evaluate the accuracy of 3D understanding as the agreement between estimated volume in mesh and ground truth mesh, measured in F1-score. The proposed evaluation is applicable to both NeRF and 3DGS.
@@
In the experiments, we demonstrate our superiority regarding 1) 3D and 2D segmentation accuracy, 2) training and rendering time, and 3) consistency across viewpoints.
In short, our contributions are:
@@
We propose to directly supervise 3D points before volume rendering to learn a language embedding field. It achieves the state-of-the-art accuracy in 3D and 2D segmentation.
We propose to transfer the language field to 3DGS. It achieves the first real-time rendering of semantics which is 27 faster than the previous fastest method.
@@ 2 Related Work
LERF (Kerr et al. 2023) pre-computes multi-scale patches to prepare multi-scale ground-truth CLIP features. Instead of using patches, LEGaussians (Shi et al. 2023) distills pixel-level CLIP features. Similarly, OpenNeRF (Engelmann et al. 2024) computes pixel-level CLIP features using OpenSeg (Ghiasi et al. 2022). As it produces mixed CLIP features to contain multiple objects into a patch, LangSplat (Qin et al. 2023) utilizes SAM (Kirillov et al. 2023) to create patches for a single object. Since studies using multi-scale feature maps select the most relevant scale, different scales can be chosen at different views with the same query, leading to multi-view inconsistent results (Kerr et al. 2023; Qin et al. 2023; Shi et al. 2023). On the other hand, we use SAM to make the field free from scales.
@@
Directly embedding high-dimensional CLIP features into each Gaussian is infeasible due to the limited GPU shared memory. Therefore, recent studies for understanding open-vocabulary 3D segmentation using 3D as a backbone address this problem in two ways. LEGaussians and LangSplat compress the CLIP features using an autoencoder (Shi et al. 2023; Qin et al. 2023). However, they need to optimize the autoencoder for each scene and endure performance degradation. Others using 3DGS without feature compression (Zhou et al. 2023; Bhalgat et al. 2024; Labe et al. 2024) utilize global memory, suffering for longer training time. Meanwhile, we effectively optimize 3DGS transferred from our learned language field, without the above trade-off.
@@ 3 Methods
Firstly, we redefine 3D segmentation and explain how to query 3D and 2D segmentation results. Subsequently, we construct an additional language field to NeRF as in LERF. Then, we introduce point-wise semantic loss to supervise embeddings of sampled ray points in the language field and describe our scale-free embeddings. Additionally, we propose to transfer our pre-trained language field into 3DGS for real-time rendering. Lastly, we introduce the evaluation protocol for 3D segmentation. Figure 2 illustrates an overview.
@@ 3.1 Preliminary: LERF
For open-vocabulary segmentation, LERF (Kerr et al. 2023) builds an additional language field on top of iNGP (Müller et al. 2022). The language field is jointly trained with the radiance field by querying point embeddings at sample points’ position and scale along each ray. The 3D point language embeddings are then accumulated via the volume rendering technique (Max 1995) to obtain rendered language embedding in 2D pixel space: . The weight of each sampled point is calculated as , where is transmittance, is the distance between adjacent samples, and is the volume density. Then, the rendered embedding is normalized to the unit sphere as in CLIP: To train the language field, LERF crops the training dataset into multi-scale patches, creating ground truth language embedding and maximizing the cosine similarity between the rendered language embedding :
@@ 3.2 Task Redefinition
Existing methods have limited 3D semantic understanding as they produce 2D masks for given texts and viewpoints. Instead, we propose 3D segmentation as a more practical problem setting: segmenting the 3D volume for a given text. While following the basic elements of language embeddings and relevancy scores, we compute relevancy scores of the language embeddings queried on 3D points instead of the ones rendered on 2D images.
where represents predefined canonical texts such as photo and image. In NeRF representation, we compute the relevancy scores on the points along the rays through pixels. For 3DGS representation, we compute the relevancy scores on the center positions of the Gaussians.
For 2D segmentation, we follow the existing works to compute the relevancy score using the rendered embeddings without scale .
The regions where the computed relevancy scores surpass the selected threshold are considered object regions.
@@ 3.3 Supervising Semantics in 3D Space
It is a reasonable choice to minimize the error between the rendered colors and the ground truth colors on 2D images for novel view synthesis. On the other hand, the similar objective for language embeddings (Eq. (1)) harms correctly understanding 3D semantics as shown in Figure 1.
To address this issue, we propose a point-wise semantic loss which uses CLIP embeddings as direct ground truth supervision for the embeddings of 3D points on the ray . Specifically, we maximize the similarity between point embeddings and the ground truth CLIP embedding :
We also apply the same approach described above for DINO regularization. We show that this improves not only 3D segmentation performance but also 2D segmentation performance thanks to a better understanding of 3D scenes.
We observe that existing works (Kerr et al. 2023; Qin et al. 2023) on multi-scale CLIP embedding fields require different optimal scales for a text query based on the viewpoint, leading to multi-view inconsistency. To address this, we use SAM (Kirillov et al. 2023) to obtain object masks and create scale-free ground truth CLIP embeddings. This ensures view-consistent language fields without the need to determine the optimal scale. We note that previous methods with SAM still model multi-scale CLIP embedding fields (Zhang, Li, and Ahuja 2024; Qin et al. 2023).
@@ 3.4 Transferring Language Field into 3DGS
A straightforward approach for the same task with 3DGS is to jointly optimize the Gaussians and their additional language embeddings, which requires vast memory consumption larger than GPU shared memory. It suffers from slow training (47) by allocating GPU global memory111Training on GPU global memory slows due to the gradient computation in the backward pass. The forward pass is still fast. (Zhou et al. 2023; Bhalgat et al. 2024), or requires an additional model for feature compression, which degrades accuracy due to compression loss (Shi et al. 2023; Qin et al. 2023). Appendix B.1 provides more details.
To tackle this problem, we propose to transfer our pre-trained language field into 3DGS, removing the need for independent optimization. We simply append the language embeddings queried at the center of Gaussians. It runs instantly (22ms) and still allows fast rendering11footnotemark: 1.
In addition, we propose to initialize the coordinates and language embeddings of 3DGS from the pre-trained radiance-language field in order to ensure the agreement between the geometry and semantics. We collect ray points and extract the top 1M points regarding density following NeRFstudio (Tancik et al. 2023). During training, we freeze the center position of the Gaussians , without densification or pruning. We then concatenate language embeddings at the center of Gaussians.
Similar to previous methods, we modify the rasterizer to use global memory instead of limited shared memory to render high-dimensional language features. The rendered language embedding is obtained by -blending:
@@
A previous work (Engelmann et al. 2024) follows the 2D segmentation protocol for 3D segmentation: mIoU on the ground truth point clouds instead of 2D pixels. It might incorrectly classify the points with zero density and correct embedding, as shown in Figure 5(a).
To address this problem, we propose to measure the F1-score between the exported mesh from the segmented results to ground truth mesh, inspired by surface reconstruction literature (Knapitsch et al. 2017; Li et al. 2023). Precision computes the ratio of correct points among the estimated mesh vertices. Recall computes the ratio of covered points among the ground truth mesh vertices. A point is considered correct or covered if there exists a point within a radius in the counterpart. We note that this protocol can be generally applied to various neural representations, such as NeRF or 3DGS.
| Method | Replica | LERF | 3D-OVS |
| LERF | 0.50 | 0.50 | 0.50 |
| OpenNeRF | 0.50 | 0.50 | 0.50 |
| Ours-NeRF | 0.55 | 0.60 | 0.55 |
| LEGaussians | 0.50 | 0.50 | 0.50 |
| LangSplat | 0.40 | 0.40 | 0.40 |
| Ours-3DGS | 0.60 | 0.60 | 0.55 |
@@ 4 Experiments
In this section, we evaluate our methods on various datasets and compare them to competitors in terms of 3D and 2D segmentation with given text queries. We choose LERF, OpenNeRF, LEGaussians, and LangSplat as competitors, which are open-sourced. We used the official code without modification. Appendix A.2 provides details of the datasets.
@@
For quantitative comparison in 3D segmentation, we export meshes using the Poisson surface reconstruction. In LERF, OpenNeRF, and Ours-NeRF, we use Nerfstudio (Tancik et al. 2023) to export meshes with 30K sampled points. In LEGaussians, LangSplat and Ours-3DGS, we use SuGaR (Guédon and Lepetit 2023) to export meshes. For 2D segmentation, we evaluate the mIoU and mAP between the ground truth mask and the predicted mask. For qualitative comparison in 3D segmentation, we use 50K points for Nerfstudio in LERF, OpenNeRF, and Ours-NeRF. We use the thresholds in Table 1 for quantitative results.
@@ 4.1 Segmentation Accuracy
| Backbone | Method | LERF | 3D-OVS | ||
| mIoU | mAP | mIoU | mAP | ||
| NeRF | LERF | 31.88 | 30.44 | 52.60 | 57.03 |
| OpenNeRF | 26.52 | 25.84 | 75.12 | 75.35 | |
| Ours-NeRF | 46.37 | 45.86 | 77.46 | 84.65 | |
| 3DGS | LEGaussians | 21.43 | 20.81 | 47.98 | 50.86 |
| LangSplat | 37.53 | 36.39 | 74.54 | 79.49 | |
| Ours-3DGS | 44.37 | 44.57 | 77.51 | 84.86 | |
2D Segmentation
In Table 3, both of our models show the highest 2D segmentation performance on LERF and 3D-OVS datasets. Also, we present qualitative results of 2D segmentation by comparing a heatmap for given text queries through 2D querying (Section 3.2). As shown in Figure 7, Ours-NeRF and Ours-3DGS show clear segmentation boundaries. LangSplat and LEGaussians occasionally fail to find the target object, as seen with waldo.
@@
| Ours-3DGS | 40+5 mins | 14.740 ms | - | - | 67.8 | ✓ | |
| Method | 3D Segmentation | 2D Segmentation | Computational Cost | |||
@@
| Ours-3DGS | 0.1354 | 44.37 | 44.57 | 5 mins | 67.8 | 1M |
| w/o NeRF init | 0.1108 | 36.14 | 36.74 | 13 mins | 24.8 | 3M |

### v3

@@ Abstract
Understanding the 3D semantics of a scene is a fundamental problem for various scenarios such as embodied agents. While NeRFs and 3DGS excel at novel-view synthesis, previous methods for understanding their semantics have been limited to incomplete 3D understanding: their segmentation results are rendered as 2D masks that do not represent the entire 3D space. To address this limitation, we redefine the problem to segment the 3D volume and propose the following methods for better 3D understanding. We directly supervise the 3D points to train the language embedding field, unlike previous methods that anchor supervision at 2D pixels. We transfer the learned language field to 3DGS, achieving the first real-time rendering speed without sacrificing training time or accuracy. Lastly, we introduce a 3D querying and evaluation protocol for assessing the reconstructed geometry and semantics together. Code, checkpoints, and annotations are available at the project page.
Project page — https://hyunji12.github.io/Open3DRF/
@@ 1 Introduction
Semantically understanding 3D space is important for various computer vision tasks. For instance, it is crucial to segment 3D objects accurately for robot manipulation (Rashid et al. 2023; Zheng et al. 2024). Recently, several works have focused on understanding 3D scenes represented by radiance fields such as Neural Radiance Fields (NeRFs) (Mildenhall et al. 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023). LERF (Kerr et al. 2023) introduces a language embedding field which is rendered on a chosen viewpoint to be queried by open vocabulary. The language field is supervised by CLIP (Radford et al. 2021) embeddings from the multi-scale patches to capture various sizes of objects. Subsequent works (Zhang, Li, and Ahuja 2024; Qin et al. 2023) incorporate SAM masks (Kirillov et al. 2023) to supervise the language field for clear segmentation boundaries.
@@
We revisit the 3D understanding of NeRFs and 3DGS in four aspects: problem setting, supervision, embeddings, and evaluation. The problem setting of previous works leads to limited 3D understanding as they merely produce rendered 2D masks for given viewpoints rather than 3D semantic volumes. On the other hand, we set the problem to segment 3D volume regarding the semantics of 3D points.
@@
The previous methods encourage the rendered embeddings, rather than the embeddings in 3D space, to match the ground truth. Furthermore, their multi-scale language embeddings require finding the optimal scale for a given viewpoint and a query text, which is prone to multi-view inconsistency. In contrast, we encourage 3D points to learn language embeddings following our revisited problem setting. Although the ground truth is still language embeddings of 2D images, changing the dimension for computing the loss greatly improves 3D understanding as shown in Figure 1. In addition, we remove multi-scale embeddings by encoding masked objects, achieving multi-view consistent rendered feature maps.
@@
Meanwhile, previous works on understanding 3DGS explicitly add language embeddings for each Gaussian and jointly train them (Qin et al. 2023; Zhou et al. 2023). However, directly appending a 512-dimensional language embedding to each Gaussian and rendering them result in out-of-memory (OOM) issues. Hence, they either 1) compress language embeddings to low-dimensional features, significantly sacrificing the accuracy (Shi et al. 2023; Qin et al. 2023), or 2) modify the rasterizer, which is slow for training (Zhou et al. 2023; Ye and Kanazawa 2023). To overcome these drawbacks, we transfer our learned language field into 3DGS to enable the real-time rendering of the accurate language field.
@@
Our final aspect is evaluation. Existing works measure mIoU in pixels between the rendered 2D masks and the 2D ground truth, or measure mIoU on ground truth point clouds ignoring the reconstructed geometry (Engelmann et al. 2024). However, 3D understanding should know the correct 3D volume of a target semantics. Therefore, the evaluation method should assess semantics and the reconstructed geometry together. Accordingly, we propose to evaluate the accuracy of 3D understanding as the agreement between estimated volume in mesh and ground truth mesh, measured in F1-score. The proposed evaluation is applicable to both NeRF and 3DGS.
@@
In the experiments, we demonstrate the superiority of our method regarding 1) 3D and rendered 2D segmentation accuracy, 2) training and rendering time, and 3) consistency across viewpoints.
In summary, our contributions are:
@@
We propose to directly supervise 3D points before volume rendering to learn a language embedding field. It achieves the state-of-the-art accuracy in 3D and rendered 2D segmentation.
We propose to transfer the language field to 3DGS. It achieves the first real-time rendering speed among open-vocabulary methods which is 28 faster than the previous fastest method.
@@ 2 Related Work
LERF (Kerr et al. 2023) pre-computes multi-scale patches to prepare multi-scale ground-truth CLIP features. Similar to LERF, FMGS (Zuo et al. 2024) pre-computes multi-scale ground-truth CLIP features and averages them to generate a low-resolution hybrid feature map for training. Instead of using patches, LEGaussians (Shi et al. 2023) distills pixel-level CLIP features. Similarly, OpenNeRF (Engelmann et al. 2024) computes pixel-level CLIP features using OpenSeg (Ghiasi et al. 2022). GOI (Qu et al. 2024) trains the codebook to compress the high-dimensional semantic features into low-dimensional vectors111At the time of this work, the code of GOI was unavailable. However, its released code showed 15 fps on LERF dataset, falling short of real-time rendering.. As it produces mixed CLIP features to contain multiple objects into a patch, LangSplat (Qin et al. 2023) utilizes SAM (Kirillov et al. 2023) to create patches for a single object. OpenMask3D (Takmaz et al. 2023) creates a bounding box from the SAM mask, and crops the image to make per-mask features. However, OpenMask3D uses bounding box, which still leads to mixed CLIP features. Since studies using multi-scale feature maps select the most relevant scale, different scales can be chosen at different views with the same query, leading to multi-view inconsistent results (Kerr et al. 2023; Qin et al. 2023; Shi et al. 2023). On the other hand, we use SAM to make the field free from scales.
@@
Directly embedding high-dimensional CLIP features into each Gaussian is infeasible due to the limited GPU shared memory. Therefore, recent studies for understanding open-vocabulary 3D segmentation using 3D as a backbone address this problem in two ways. LEGaussians and LangSplat compress the CLIP features using an autoencoder (Shi et al. 2023; Qin et al. 2023). However, they need to optimize the autoencoder for each scene and endure performance degradation. Others using 3DGS without feature compression (Zhou et al. 2023; Labe et al. 2024) utilize global memory, suffering for longer training time. Meanwhile, we effectively optimize 3DGS transferred from our learned language field, without the above trade-off.
@@ 3 Methods
First, we provide preliminary on LERF (Kerr et al. 2023). Then we redefine 3D segmentation of radiance fields as computing 3D relevancy scores. Accordingly, we introduce point-wise semantic loss to supervise ray points in 3D space by our scale-free embeddings. In addition, we propose to transfer our learned language field into 3DGS for real-time rendering. Lastly, we introduce the evaluation protocol for 3D segmentation. Figure 2 illustrates an overview.
@@ 3.1 Preliminary: LERF
For open-vocabulary segmentation, LERF builds an additional language field on top of iNGP (Müller et al. 2022). The language field is jointly trained with the radiance field by querying point embeddings at sample points’ position and scale along each ray. The 3D point language embeddings are then accumulated via the volume rendering (Max 1995) to obtain rendered language embedding in 2D pixel space: . The weight of each sampled point is calculated as , where is transmittance, is the distance between adjacent samples, and is the volume density. Then, the rendered embedding is normalized to the unit sphere as in CLIP: To train the language field, LERF crops the training dataset into multi-scale patches, creating ground truth language embedding and maximizing the cosine similarity between the rendered language embedding :
@@
For rendered 2D segmentation, LERF computes the relevancy score using the rendered embeddings and obtains the 2D mask:
@@ 3.2 Task Redefinition
Existing methods provide limited 3D semantic understanding for given texts via rendered 2D masks on viewpoints that do not directly represent the entire 3D space. Instead, we propose 3D segmentation as a more practical problem setting: segmenting the 3D volume for given texts. While following the basic elements of language embeddings and relevancy scores, we compute relevancy scores of the language embeddings queried on 3D points instead of the ones rendered on 2D images:
where represents predefined canonical texts such as photo and image. In NeRFs, we compute the relevancy scores on the points along the rays through pixels. For 3DGS, we compute the relevancy scores on the center positions of the Gaussians. The regions where the computed relevancy scores surpass the selected threshold are considered object regions.
@@ 3.3 Supervising Semantics in 3D Space
It is a reasonable choice to minimize the error between the rendered colors and the ground truth colors on 2D images for novel view synthesis. In contrast, the similar objective for language embeddings (Eq. (1)) harms correctly understanding 3D semantics as shown in Figure 1.
To address this issue, we propose a point-wise semantic loss which uses CLIP embeddings as direct ground truth for the embeddings of 3D points on the ray . Specifically, we maximize the similarity between point embeddings and the ground truth CLIP embedding :
We also apply the same approach described above for DINO regularization. We show that point-wise semantic loss improves not only 3D segmentation but also rendered 2D segmentation thanks to a better understanding of 3D semantics.
Notably, multi-scale approaches (Kerr et al. 2023; Qin et al. 2023) lead to multi-view inconsistency because these approaches have language embedding per scale and different views may have different optimal scales. To address this, we use SAM (Kirillov et al. 2023) to obtain object masks and create scale-free ground truth CLIP embeddings by masking out their background and cropping tightly to the masks. It ensures view-consistent language fields without the need to determine the optimal scale. We note that previous methods with SAM still model multi-scale CLIP embedding fields (Zhang, Li, and Ahuja 2024; Qin et al. 2023).
@@ 3.4 Transferring Language Field into 3DGS
A straightforward approach for the same task with 3DGS is to jointly optimize the Gaussians and their additional language embeddings, which requires vast memory consumption. It suffers from slow training222Training on GPU global memory is slow due to the gradient computation in the backward pass. The forward pass is still fast. (Zhou et al. 2023), or requires an additional model for feature compression, which degrades accuracy due to compression loss (Shi et al. 2023; Qin et al. 2023). Appendix provides more details.
To tackle this problem, we propose to simply bake our learned language field into 3DGS by querying the language embedding at the center of the Gaussians. The queried embeddings are frozen. It runs instantly (22ms) and still allows fast rendering22footnotemark: 2.
In addition, we propose to initialize the center coordinates of 3DGS from the learned radiance field. We collect ray points and extract the top 1M points regarding density following NeRFstudio (Tancik et al. 2023). Unlike previous methods (Niemeyer et al. 2024), we implement this approach to align the geometry of 3DGS with the language field. Therefore, we freeze the center position of the Gaussians , without densification or pruning during training. We then transfer language embeddings at the center of Gaussians.
Similar to previous methods, we modify the rasterizer to render high-dimensional language features (see Appendix). The rendered language embedding is obtained by -blending:
@@
A previous work (Engelmann et al. 2024) measures the mIoU between the prediction and annotation on the ground truth point cloud. It might incorrectly classify the points with zero density and a correct embedding, shown in Figure 5-(a).
To address this problem, we propose to measure the F1-score between the exported mesh from the segmented results to ground truth mesh, inspired by surface reconstruction literature (Knapitsch et al. 2017; Li et al. 2023). Precision computes the ratio of correct volume among the estimated volume. Recall computes the ratio of covered volume among the GT volume. As it is impractical to compute the volume, we approximate the volume by a regular grid points in the estimated and GT meshes. A point is considered correct or covered if there exists a point within a radius in the counterpart. We note that this protocol can be generally applied to various neural representations, such as NeRF or 3DGS.
@@ 4 Experiments
In this section, we evaluate our methods on various datasets and compare them to competitors in terms of 3D and rendered 2D segmentation with given text queries. We choose LERF, OpenNeRF, LEGaussians, and LangSplat as competitors, which are open-sourced. We re-evalute the competitors using the official code333The official implementation of LangSplat includes evaluation views during training, while we exclude them for fairness, following LERF and OpenNeRF.. Appendix provides details of the datasets.
@@
For quantitative comparison in 3D segmentation, we export meshes using the Poisson surface reconstruction. In LERF, OpenNeRF, and Ours-NeRF, we use Nerfstudio (Tancik et al. 2023) to export meshes with 30K sampled points. In LEGaussians, LangSplat and Ours-3DGS, we use SuGaR (Guédon and Lepetit 2023) to export meshes. For rendered 2D segmentation, we evaluate the mIoU and mAP between the ground truth mask and the predicted mask. For qualitative comparison in 3D segmentation, we use 50K points for Nerfstudio in LERF, OpenNeRF, and Ours-NeRF.
@@ 4.1 Segmentation Accuracy
Method | LERF | 3D-OVS | ||
mIoU | mAP | mIoU | mAP | |
LERF | 31.88 | 30.44 | 52.60 | 57.03 |
OpenNeRF | 26.52 | 25.84 | 75.12 | 75.35 |
Ours-NeRF | 46.37 | 45.86 | 77.46 | 84.65 |
LEGaussians | 21.43 | 20.81 | 47.98 | 50.86 |
LangSplat33footnotemark: 3 | 37.53 | 36.39 | 74.54 | 79.49 |
Ours-3DGS | 44.37 | 44.57 | 77.51 | 84.86 |
Rendered 2D Segmentation
In Table 2, both of our models show the highest rendered 2D segmentation performance on LERF and 3D-OVS datasets. Also, we present qualitative results of rendered 2D segmentation by comparing a heatmap for given text queries through 2D querying (Section 3.2). As shown in Figure 7, Ours-NeRF and Ours-3DGS show clear segmentation boundaries. LangSplat and LEGaussians occasionally fail to find the target object, as seen with waldo.
@@
Ours-3DGS | 40+5 mins | 14.257 ms | - | - | 70.1 | ✓ |
Method | Replica | LERF | Computational Cost | |||
@@
Ours-3DGS | 0.1354 | 44.37 | 44.57 | 5 mins | 70.1 | 1M |
w/o NeRF init | 0.1108 | 36.14 | 36.74 | 13 mins | 41.2 | 2M |